OpenStudy (anonymous):

Can someone please help me? I need to know the answer to this (simplified answer). Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

i think it is b2-4ac

OpenStudy (anonymous):

I am prett sure that is the answer to the first part but you are suppose to use that to find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

And that is what I don't know the answer to? Can you help?

OpenStudy (anonymous):

if you want to find the value of x, \[x = \frac{ -b \pm \sqrt{b ^{2}- 4ac} }{ 2a }\] a= coefficient of the x^2 term b=coefficient of the x term c=constant

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you are welcome!!

OpenStudy (anonymous):

:)